### v1

# eKalibr-Stereo: Continuous-Time Spatiotemporal Calibration for Event-Based Stereo Visual Systems

Abstract
The bioinspired event camera, distinguished by its exceptional temporal resolution, high dynamic range, and low power consumption, has been extensively studied in recent years for motion estimation, robotic perception, and object detection. In ego-motion estimation, the stereo event camera setup is commonly adopted due to its direct scale perception and depth recovery. For optimal stereo visual fusion, accurate spatiotemporal (extrinsic and temporal) calibration is required. Considering that few stereo visual calibrators orienting to event cameras exist, based on our previous work eKalibr (an event camera intrinsic calibrator), we propose eKalibr-Stereo for accurate spatiotemporal calibration of event-based stereo visual systems. To improve the continuity of grid pattern tracking, building upon the grid pattern recognition method in eKalibr, an additional motion prior-based tracking module is designed in eKalibr-Stereo to track incomplete grid patterns. Based on tracked grid patterns, a two-step initialization procedure is performed to recover initial guesses of piece-wise B-splines and spatiotemporal parameters, followed by a continuous-time batch bundle adjustment to refine the initialized states to optimal ones. The results of extensive real-world experiments show that eKalibr-Stereo can achieve accurate event-based stereo spatiotemporal calibration. The implementation of eKalibr-Stereo is open-sourced at (https://github.com/Unsigned-Long/eKalibr) to benefit the research community.
Index Terms:
Stereo event camera, spatiotemporal calibration, continuous-time optimization, event-based circle grid recognition
I Introduction and Related Works
Bioinspired event cameras have attracted considerable research interest in recent years, due to their advantages of low sensing latency and high dynamic range over conventional standard (frame-based) cameras [1]. The ego-motion estimation in high-dynamic-range and high-speed scenarios is one of applications of the event camera, where a stereo camera setup is commonly employed for direct scale recovery [2, 3, 4]. For such an event-based stereo visual sensor suite, accurate spatiotemporal calibration is required to determine extrinsics and time offset between cameras for subsequent data fusion.
Figure 1: The runtime visualization of eKalibr-Stereo. eKalibr-Stereo tracks grid patterns using raw events of stereo event cameras and estimates spatiotemporal (extrinsic and temporal) parameters using continuous-time-based batch optimization.
Stereo visual spatiotemporal calibration typically consists of two sub-modules: ($i$) correspondence construction (front end) and ($ii$) spatiotemporal optimization (back end). In the front end, artificial visual targets, such as checkerboards [5], April Tags [6], and ChArUco board [7], are commonly employed to construct accurate 3D-2D correspondences with real-world geometric scale through pattern recognition. While a substantial number of target pattern recognition methods [5, 8, 9] oriented to standard cameras have been proposed, they are not applicable to event cameras, which output asynchronous event stream rather than conventional intensity images. To recognize target patterns from raw events, early works [10, 11, 12] generally rely on blinking light emitting diode (LED) grid boards. Although target patterns can be accurately extracted, requiring additional LED boards introduces inconvenience. Meanwhile, these methods typically require the event camera to remain stationary, making them unsuitable for multi-camera spatiotemporal calibration that necessitates motion excitation [13]. To address this, subsequent methods [14, 15] have proposed an alternative approach, namely reconstructing intensity images from raw events using event-based image reconstruction methods (such as E2VID[16] and Spade-E2VID [17]) first, followed by conventional image-based pattern recognition methods. Although reconstructed images exhibit high consistency, substantial noise within the images could lead to imprecise pattern extraction, which further affects calibration accuracy. Considering these, some event-based pattern recognition methods have been proposed recently, aiming to extract target patterns from dynamically acquired raw events directly. Based on density-based spatial clustering (DBSCAN), Huang et al. [18] cluster events accumulated in a time window and fit circles to extract centers of circles on a circle grid board. Similarly but performing DBSCAN in spatiotemporal domain, Salah et al. [19] cluster events generated from a circle grid board, and then fit 3D cylinders for circle center determination. Focusing on event-frame stereo camera calibration, Wang et al. [20] designed a novel circle grid board with cross points, and match candidate event regions of opposite polarity for 3D oblique elliptical cylinder fitting. In our previous work eKalibr [21], an event-only target pattern recognition method oriented to the circle grid board is proposed for intrinsic calibration, which cluster events and match clustering pair based on normal flow estimation, and fit 3D elliptical cylinders for circle center extraction.
In terms of the back end of the stereo visual calibration, namely spatiotemporal optimization, event-based and frame-based calibrations share the same algorithmic framework, aiming to estimate spatiotemporal parameters using extracted visual target patterns. In general, spatiotemporal optimization can be categorized into discrete-time-based and continuous-time-based ones. Discrete-time-based methods represent states using discrete estimates that are temporally coupled to measurements. Based on the extended Kalman filter (EKF), Mirzaei et al. [22] proposed a visual-inertial extrinsic calibration method to determine the transformation between a standard camera and an inertial measurement unit (IMU). Similarly, Hartzer et al. [23] presented an EKF-based online visual-inertial extrinsic calibration method. Yang et al. [24] designed a sliding-window-based visual-inertial state estimator, supporting online camera-IMU extrinsic calibration. Different from the discrete-time-based methods, continuous-time-based ones represent time-varying states using time-continuous functions (such as B-splines), enabling state querying at any time instance, and thus are more suitable for temporal calibration. The well-known Kalibr [25] proposed by Furgale et al. is the first continuous-time-based calibration framework, which employs B-splines for state representation and supports both extrinsic and temporal calibration for visual-inertial, multi-IMU, and multi-camera sensor suites. Kalibr is then extended by Huai et al. [26] to support the rolling shutter cameras for readout time calibration. In addition to vision-related calibration, the continuous-time state representation has also been widely employed in other multi-sensor calibration, such as LiDAR-IMU [27] and radar-IMU [28] calibration.
In this article, focusing on event-based stereo visual systems, we present a continuous-time-based spatiotemporal calibration method, named eKalibr-Stereo, to accurately estimate the extrinsics and time offset between event cameras. Building upon our previous work eKalibr [21], eKalibr-Stereo tracks continuous circle grid patterns (complete and incomplete ones) from raw events for 3D-2D correspondence construction. Given the high non-linearity of continuous-time optimization, a two-stage initialization procedure is first conducted to recover the initials of states, which are then iteratively refined to optimal ones using a continuous-time-based batch bundle adjustment. eKalibr-Stereo makes the following (potential) contributions:
  1. 1.
    We proposed a continuous-time-based spatial and temporal calibrator for event-based stereo visual systems, which could accurately determine both extrinsics and time offset of a stereo event camera system.
  2. 2.
    We designed a motion-prior-aided tracking module for incomplete grid pattern identification, to maximize the continuity of pattern tracking, facilitating the final spatiotemporal optimization.
  3. 3.
    Sufficient real-world experiments were conducted to comprehensively evaluate the proposed eKalibr-Stereo. Both the dataset and code implementation are open-sourced, to benefit the robotic community if possible.
Note that the proposed eKalibr-Stereo supports one-shot event-based multi-camera spatiotemporal calibration (an arbitrary number of event cameras). To enhance clarity, this article only considers the minimal stereo event camera configuration, as it’s the most critical sensor setup for facilitating multi-camera calibration.
II Preliminaries
This section presents notations and definitions utilized in this article. The involved camera intrinsic model and B-spline-based time-varying state representation are also introduced for a self-contained exposition of this work.
II-A Notations and Definitions
Given a raw event $\boldsymbol{\mathrm{e}}$ generated by the event camera, we use $\tau\in\mathbb{R}$, $\boldsymbol{\mathrm{x}}\in\mathbb{Z}^{2}$, and $p\in\{{\text{-}}1,{\text{+}}1\}$ to represent its timestamp, pixel position, and polarity respectively, i.e., $\boldsymbol{\mathrm{e}}\triangleq\{\tau,\boldsymbol{\mathrm{x}},p\}$. The camera frame and world frame (defined by the circle grid board) are represented as $\underrightarrow{\mathcal{F}}_{c}$ and $\underrightarrow{\mathcal{F}}_{w}$, respectively. The transformation from $\underrightarrow{\mathcal{F}}_{c}$ to $\underrightarrow{\mathcal{F}}_{w}$ are parameterized as the Euclidean matrix ${\boldsymbol{\mathrm{T}}_{c}^{w}}\in\mathrm{SE(3)}$, which is defined as:
${\boldsymbol{\mathrm{T}}_{c}^{w}}\triangleq\begin{bmatrix}{\boldsymbol{\mathrm{R}}_{c}^{w}}&{\boldsymbol{\mathrm{p}}_{c}^{w}}\\ \boldsymbol{\mathrm{0}}_{1\times 3}&1\end{bmatrix}$ (1)
where ${\boldsymbol{\mathrm{R}}_{c}^{w}}\in\mathrm{SO(3)}$ and ${\boldsymbol{\mathrm{p}}_{c}^{w}}\in\mathbb{R}^{3}$ are the rotation matrix and translation vector, respectively. Finally, we use $\hat{(\cdot)}$ and $\tilde{(\cdot)}$ to represent the state estimates and noisy quantities (e.g., the generated raw events and extracted grid patterns), respectively.
II-B Camera Intrinsic Model
The camera intrinsic model characterizes the visual projection process whereby 3D points in the camera coordinate frame are geometrically mapped onto the image plane to derive corresponding 2D pixels. Adhering to our previously proposed eKalibr [21], the intrinsic camera model comprising the pinhole projection model [29] and radial-tangential distortion model [30] are employed in this work, which can be expressed as:
$\boldsymbol{\mathrm{x}}_{p}=\pi\left(\boldsymbol{\mathrm{p}}^{c},\mathcal{X}_{\mathrm{intr}}\right)\triangleq\boldsymbol{\mathrm{K}}\left(\mathcal{X}_{\mathrm{proj}}\right)\cdot\boldsymbol{\mathrm{d}}\left(\boldsymbol{\mathrm{p}}^{c},\mathcal{X}_{\mathrm{dist}}\right)$ (2)
with
$\begin{gathered}\mathcal{X}_{\mathrm{intr}}\triangleq\mathcal{X}_{\mathrm{proj}}\cup\mathcal{X}_{\mathrm{dist}}\\ \mathcal{X}_{\mathrm{proj}}\triangleq\left\{f_{x},f_{y},c_{x},c_{y}\right\},\;\mathcal{X}_{\mathrm{dist}}\triangleq\left\{k_{1},k_{2},p_{1},p_{2}\right\}\end{gathered}$ (3)
where $\boldsymbol{\mathrm{d}}:\mathbb{R}^{3}\mapsto\mathbb{R}^{3}$ represents the distortion function distorting normalized image coordinates using distortion parameters $\mathcal{X}_{\mathrm{dist}}$; $\boldsymbol{\mathrm{K}}\in\mathbb{R}^{2\times 3}$ denotes the intrinsic matrix organized by projection parameters $\mathcal{X}_{\mathrm{proj}}$; $\pi:\mathbb{R}^{3}\mapsto\mathbb{R}^{2}$ is the projection function projecting 3D point $\boldsymbol{\mathrm{p}}^{c}$ onto the image plane as 2D point $\boldsymbol{\mathrm{x}}_{p}$; $\mathcal{X}_{\mathrm{intr}}$ represents the intrinsic parameters comprising $\mathcal{X}_{\mathrm{proj}}$ and $\mathcal{X}_{\mathrm{dist}}$, which can be pre-calibrated using eKalibr.
II-C Continuous-Time State Representation
To efficiently fuse asynchronous data for multi-sensor spatiotemporal determination, especially for time offset calibration, the continuous-time state representation is employed in this work to represent the time-varying rotation and position of event cameras. Compared with the conventional discrete-time representation generally maintaining discrete states at measurement times, the continuous-time representation models time-varying states using time-continuous functions, such as Gaussian process regression [31], hierarchical wavelets [32], and B-splines [33], enabling state querying at arbitrary time. In this work, the uniform B-spline is utilized for continuous-time state representation, which inherently possesses sparsity due to its local controllability, allowing computation acceleration in optimization [13].
The uniform B-spline is characterized by the spline order, a temporally uniformly distributed control point sequence, and a constant time distance between neighbor control points. Specifically, given a series of translational control points:
$\begin{gathered}\mathcal{X}_{\mathrm{pos}}\triangleq\left\{\boldsymbol{\mathrm{p}}_{i},\tau_{i}\mid\boldsymbol{\mathrm{p}}_{i}\in\mathbb{R}^{3},\tau_{i}\in\mathbb{R}\right\}\\ \mathrm{s.t.}\;\;\tau_{i{\text{+}}1}-\tau_{i}\equiv\Delta\tau_{\mathrm{pos}}\end{gathered}$ (4)
the position $\boldsymbol{\mathrm{p}}(\tau)$ at time $\tau\in[\tau_{i},\tau_{i{\text{+}}1})$ of a $k$-order uniform B-spline can be computed as follows:
$\begin{gathered}\boldsymbol{\mathrm{p}}(\tau)=\boldsymbol{\mathrm{p}}_{i}+\sum_{j=1}^{k{\text{+}}1}\lambda_{j}(u)\cdot\left(\boldsymbol{\mathrm{p}}_{i{\text{+}}j}-\boldsymbol{\mathrm{p}}_{i{\text{+}}j{\text{-}}1}\right)\\ \mathrm{s.t.}\;\;u=\frac{\tau-\tau_{i}}{\Delta\tau_{\mathrm{pos}}}\end{gathered}$ (5)
where $\lambda_{j}(\cdot)$ denotes the $j$-th element of vector $\boldsymbol{\mathrm{\lambda}}(u)$ obtained from the order-determined cumulative matrix and $u$ [13]. In this work, the cubic uniform B-spline ($k=4$) is employed.
The B-spline representation of time-varying rotation has similar forms with (5) by replacing vector addition in $\mathbb{R}^{3}$ with group multiplication in $\mathrm{SO(3)}$. The key distinction resides in the scalar multiplication operated within the Lie algebra $\mathfrak{so}(3)$, rather than on the Lie group manifold, to ensure closedness [34]. Specifically, given a series of rotational control points:
$\begin{gathered}\mathcal{X}_{\mathrm{rot}}\triangleq\left\{\boldsymbol{\mathrm{R}}_{i},\tau_{i}\mid\boldsymbol{\mathrm{R}}_{i}\in\mathrm{SO(3)},\tau_{i}\in\mathbb{R}\right\}\\ \mathrm{s.t.}\;\tau_{i{\text{+}}1}-\tau_{i}\equiv\Delta\tau_{\mathrm{rot}}\end{gathered}$ (6)
the rotation $\boldsymbol{\mathrm{R}}(\tau)$ at time $\tau\in[\tau_{i},\tau_{i{\text{+}}1})$ of a $k$-order uniform B-spline can be computed as follows:
$\begin{gathered}\boldsymbol{\mathrm{R}}(\tau)=\boldsymbol{\mathrm{R}}_{i}\cdot\prod_{j=1}^{k{\text{+}}1}\mathrm{Exp}\left(\lambda_{j}(u)\cdot\mathrm{Log}\left(\boldsymbol{\mathrm{R}}_{i{\text{+}}j{\text{-}}1}^{\top}\cdot\boldsymbol{\mathrm{R}}_{i{\text{+}}j}\right)\right)\\ \mathrm{s.t.}\;\;u=\frac{\tau-\tau_{i}}{\Delta\tau_{\mathrm{rot}}}\end{gathered}$ (7)
where $\mathrm{Exp}(\cdot)$ maps elements in the Lie algebra to the associated Lie group, and $\mathrm{Log}(\cdot)$ is its inverse operation.
III Methodology
This section details the proposed event-based continuous-time stereo visual spatiotemporal calibrator eKalibr-Stereo.
III-A System Overview
Figure 2: Illustration of the pipeline of the proposed event-based stereo visual spatiotemporal calibration method. A detailed description of the pipeline is provided in Section III-A.
The comprehensive framework of the proposed event-based stereo visual calibrator is illustrated in Fig. 2. Given raw asynchronous event streams from the stereo camera rig, we first perform normal flow estimation and ellipse fitting for each event camera, to track both complete and incomplete circle grid patterns. The event camera that tracks the grid pattern the most among two cameras would be treated as the reference (primary) camera (denoted as $\underrightarrow{\mathcal{F}}_{c_{r}}$), while the other would be assigned as the target (secondary) camera (denoted as $\underrightarrow{\mathcal{F}}_{c_{t}}$). Subsequently, for each tracked grid pattern of the reference camera, PnP [35] would be employed to estimate the camera pose based on 3D-2D correspondences. Discrete poses of the reference camera are then utilized to recover a continuous-time trajectory. The spatiotemporal parameters, i.e., extrinsics and time offset, would be also initialized based on continuous-time hand-eye alignment. Finally, a continuous-time-based visual bundle adjustment would be performed to refine all states in the estimator to the global optimal ones.
The state vector of the system can be described as follows:
$\mathcal{X}\triangleq\left\{\mathcal{X}_{\mathrm{pos}},\mathcal{X}_{\mathrm{rot}},{\boldsymbol{\mathrm{R}}_{c_{t}}^{c_{r}}},{\boldsymbol{\mathrm{p}}_{c_{t}}^{c_{r}}},{\tau_{c_{t}}^{c_{r}}}\right\}$ (8)
where $\mathcal{X}_{\mathrm{pos}}$ and $\mathcal{X}_{\mathrm{rot}}$ are translational and rotational control points defined in (4) and (6), respectively; ${\boldsymbol{\mathrm{R}}_{c_{t}}^{c_{r}}}$ and ${\boldsymbol{\mathrm{p}}_{c_{t}}^{c_{r}}}$ denote the extrinsic rotation and translation from $\underrightarrow{\mathcal{F}}_{c_{t}}$ to $\underrightarrow{\mathcal{F}}_{c_{r}}$; ${\tau_{c_{t}}^{c_{r}}}$ represents the time offset between two cameras, i.e., temporal transformation $\tau_{c_{r}}=\tau_{c_{t}}+{\tau_{c_{t}}^{c_{r}}}$ holds. The extrinsics and time offset are exactly the spatiotemporal parameters eKalibr-Stereo calibrates.
III-B Event-Based Circle Grid Tracking
Given generated raw event streams, we first employ the event-based circle grid pattern recognition algorithm [21] proposed in our previous work eKalibr to extract complete grid patterns for each camera. As described in [21], we first perform event-based normal flow estimation [36] on the surface of active event (SAE) [37] and homopolarly cluster inlier events for cluster matching. Spatiotemporal ellipses would then be estimated for each matched cluster pair for center determination of the grid circle. Finally, temporally synchronized centers would be organized as ordered grid patterns. Note that although both the asymmetric and symmetric circle grids are supported in eKalibr, the asymmetric circle grid is utilized in this work as the visual target, as it does not exhibit 180-degree ambiguity [38].
Figure 3: Schematic of incomplete grid pattern tracking. When ellipses corresponding to all gird circles are fitted (subfigure A), a complete pattern (subfigure B) can be extracted. However, due to oblique perspective (subfigure C), noisy events (subfigure D), or the grid being outside the sensing range (subfigure E), not all ellipses can be identified, resulting in incomplete pattern.
Algorithm 1 Incomplete Grid Pattern Tracking
1:Input: Extracted complete grid patterns $\mathcal{P_{\mathrm{cmp}}}$ and unorganized fitted ellipses $\mathcal{E}$ for incomplete pattern tracking.
2:Output: Temporally-ordered patterns $\mathcal{P}$ including both complete grids $\mathcal{P_{\mathrm{cmp}}}$ and incomplete ones $\mathcal{P_{\mathrm{incmp}}}$.
3:Initialize grid pattern set $\mathcal{P}\leftarrow\mathcal{P_{\mathrm{cmp}}}$, traverse order $i=2$.
4:repeat
5:     for $\left(\mathcal{G}^{k{\text{-}}1},\tau^{k{\text{-}}1}\right),\left(\mathcal{G}^{k},\tau^{k}\right),\left(\mathcal{G}^{k{\text{+}}1},\tau^{k{\text{+}}1}\right)\in\mathcal{P}$ do
6:         for center $\boldsymbol{\mathrm{p}}_{j}$ tracked three times $\boldsymbol{\mathrm{p}}_{j}^{k{\text{-}}1},\boldsymbol{\mathrm{p}}_{j}^{k},\boldsymbol{\mathrm{p}}_{j}^{k{\text{+}}1}$ do
7:              Predict $\hat{\boldsymbol{\mathrm{p}}}_{j}^{k{\text{+}}i}$ at time $\tau^{k{\text{+}}i}$ using (9), associate
8:              its nearest ellipse center $\hat{\boldsymbol{\mathrm{p}}}_{j}^{k{\text{+}}i}\simeq\boldsymbol{\mathrm{c}}_{n}^{k{\text{+}}i}\in\mathcal{E}^{k{\text{+}}i}$
9:              if $\|\hat{\boldsymbol{\mathrm{p}}}_{j}^{k{\text{+}}i}-\boldsymbol{\mathrm{c}}_{n}^{k{\text{+}}i}\|_{2}\leq d_{\mathrm{thd}}$, then $\mathcal{G}^{k{\text{+}}i}\leftarrow\boldsymbol{\mathrm{c}}_{j}^{k{\text{+}}i}$.
10:         end for
11:         Store $\mathcal{P}\leftarrow\mathcal{G}^{k{\text{+}}i}$ if tracked enough points in $\mathcal{G}^{k{\text{+}}i}$.
12:     end for
13:     Reverse order $i\leftarrow\left(-1\right)\times i$.
14:until no additional incomplete pattern tracked in last loop.
15:Note: The three-point Lagrange polynomial is defined as:
$\small\boldsymbol{\mathrm{p}}(\tau)\leftarrow L_{3}(\tau)\triangleq\sum_{k=0}^{2}\left(\boldsymbol{\mathrm{p}}^{k}\cdot\left(\prod_{l=0,l\neq k}^{2}\frac{\tau-\tau^{l}}{\tau^{k}-\tau^{l}}\right)\right).$ (9)
In addition to the aforementioned grid pattern recognition algorithm [21], this work incorporates an additional module specifically designed to track incomplete grid patterns (see Fig. 3), to improve continuity of grid tracking111 Continuity of grid tracking: the motion-based spatiotemporal calibration determines parameters based on the rigid-body constraint under continuous motion, thereby requiring motion state estimation from continuous tracking of the grid. . Specifically, leveraging the prior knowledge of motion continuity, we construct a three-point Lagrange polynomial [39] for each grid circle that had been continuously tracked three times, and then predict its position in the subsequent SAE map. When the predicted point exhibits sufficient proximity to its nearest ellipse center extracted from the subsequent SAE, we designate the newly extracted ellipse center as the corresponding position of the grid circle in the subsequent SAE. Once enough predicted grid circles were associated with ellipse centers in the subsequent SAE map, we organized a new incomplete tracked grid pattern. Note that, to ensure maximal tracking continuity, we would iteratively perform alternating forward and backward tracking of incomplete grid patterns until no additional ones can be tracked. The detailed process is summarized in Algorithm 1. For notational convenience, we denote all tracked grid patterns as:
$\mathcal{P}\triangleq\left\{\left(\mathcal{G}^{k},\tau^{k}\right)\right\}\;\mathrm{s.t.}\;\;\mathcal{G}^{k}\triangleq\left\{\left.\left(\boldsymbol{\mathrm{x}}^{k}_{j},\boldsymbol{\mathrm{p}}^{w}_{j}\right)\right|\boldsymbol{\mathrm{x}}^{k}_{j}\in\mathbb{R}^{2},\boldsymbol{\mathrm{p}}^{w}_{j}\in\mathbb{R}^{3}\right\}$ (10)
where $\mathcal{G}^{k}$ denotes the $k$-th tracked grid pattern at time $\tau^{k}$, storing tracked 2D ellipse centers $\left\{\boldsymbol{\mathrm{x}}^{k}_{j}\right\}$ and their associated 3D grid circle centers $\left\{\boldsymbol{\mathrm{p}}^{w}_{j}\right\}$ on the board. The event camera that tracks the grid pattern the most among two cameras is treated as the reference camera (denote its tracked pattern set as $\mathcal{P}_{\mathrm{ref}}$), while the other is assigned as the target camera (denote its pattern set as $\mathcal{P}_{\mathrm{tar}}$).
III-C Two-Stage State Initialization
Considering the high non-linearity of continuous-time optimization, an efficient two-stage initialization procedure is designed to orderly recover initial guesses of all parameters in the estimator.
III-C1 Continuous-Time Trajectory Initialization
We first perform PnP [35] of each tracked grid pattern for both cameras to estimate discrete camera pose sequence, i.e., recover $\{{\boldsymbol{\mathrm{T}}_{c_{r}^{k}}^{w}}\}$ using $\mathcal{P}_{\mathrm{ref}}$ and $\{{\boldsymbol{\mathrm{T}}_{c_{t}^{k}}^{w}}\}$ using $\mathcal{P}_{\mathrm{tar}}$. Subsequently, we segment poses of the reference camera into multiple sections and then construct piece-wise pose B-splines. Only those neighbor poses with ($i$) time distances smaller than $\Delta\tau_{\mathrm{thd}}$ and that ($ii$) appear continuously more than $\mathcal{N}_{\mathrm{thd}}$ times, would be considered for B-spline trajectory construction. Such a strategy is designed to ensure high-precision continuous-time trajectory initialization using sufficient discrete poses. After segmentation, piece-wise pose B-splines could be initialized by solving the following nonlinear least-squares problem:
$\small\hat{\mathcal{X}}_{\mathrm{rot}},\hat{\mathcal{X}}_{\mathrm{pos}}\leftarrow\arg\min\sum_{k=0}^{n}\left\|\mathrm{Log}\left({\hat{\boldsymbol{\mathrm{T}}}_{c_{r}}^{w}}(\tau_{r}^{k})\cdot\left({\tilde{\boldsymbol{\mathrm{T}}}_{c_{r}^{k}}^{w}}\right)^{-1}\right)\right\|^{2}$ (11)
where ${\tilde{\boldsymbol{\mathrm{T}}}_{c_{r}^{k}}^{w}}$ denotes the estimated $k$-th pose of $\underrightarrow{\mathcal{F}}_{c_{r}}$ using PnP, stamped as time $\tau_{r}^{k}$ by the clock of the reference camera; ${\hat{\boldsymbol{\mathrm{T}}}_{c_{r}}^{w}}(\tau)$ represents the pose at time $\tau$, queried from the corresponding continuous-time trajectory using (5) for ${\hat{\boldsymbol{\mathrm{p}}}_{c_{r}}^{w}}(\cdot)$ and (7) for ${\hat{\boldsymbol{\mathrm{R}}}_{c_{r}}^{w}}(\cdot)$, which exactly involves the optimization of control points of B-splines. Note that for notational simplicity, the B-spline index is omitted in (11) when performing pose querying. Readers should be aware that the pose is queried from the B-spline temporally associated with it (i.e., the timestamp of the pose to query falls in the time interval of associated B-spline).
III-C2 Spatiotemporal Initialization
After the continuous-time trajectories of the reference camera are initialized, we employ the continuous-time hand-eye alignment to initialize the extrinsics and time offset of the target camera with respect to the reference camera. This could be achieved by solving the following least-squares problem:
$\small{\hat{\boldsymbol{\mathrm{R}}}_{c_{t}}^{c_{r}}},{\hat{\boldsymbol{\mathrm{p}}}_{c_{t}}^{c_{r}}},{\hat{\tau}_{c_{t}}^{c_{r}}}\!\leftarrow\!\arg\min\!\sum_{k=0}^{n}\left\|\mathrm{Log}\!\left({\hat{\boldsymbol{\mathrm{T}}}_{c_{t}^{k{\text{+}}1}}^{c_{t}^{k}}}\!\cdot\!\left({\tilde{\boldsymbol{\mathrm{T}}}_{c_{t}^{k{\text{+}}1}}^{w}}\!\right)^{-1}\!\cdot\!{\tilde{\boldsymbol{\mathrm{T}}}_{c_{t}^{k}}^{w}}\right)\right\|^{2}$ (12)
with
$\small{\hat{\boldsymbol{\mathrm{T}}}_{c_{t}^{k{\text{+}}1}}^{c_{t}^{k}}}=\left({\hat{\boldsymbol{\mathrm{T}}}_{c_{t}}^{c_{r}}}\right)^{-1}\!\cdot\left({\boldsymbol{\mathrm{T}}_{c_{r}}^{w}}(\tau_{t}^{k}+{\hat{\tau}_{c_{t}}^{c_{r}}})\right)^{-1}\!\cdot{\boldsymbol{\mathrm{T}}_{c_{r}}^{w}}(\tau_{t}^{k{\text{+}}1}+{\hat{\tau}_{c_{t}}^{c_{r}}})\cdot{\hat{\boldsymbol{\mathrm{T}}}_{c_{t}}^{c_{r}}}$ (13)
where ${\tilde{\boldsymbol{\mathrm{T}}}_{c_{t}^{k}}^{w}}$ and ${\tilde{\boldsymbol{\mathrm{T}}}_{c_{t}^{k{\text{+}}1}}^{w}}$ are two consecutive poses of the target camera obtained by PnP, stamped as $\tau_{t}^{k}$ and $\tau_{t}^{k{\text{+}}1}$ by the clock of the target camera; ${\hat{\boldsymbol{\mathrm{T}}}_{c_{t}^{k{\text{+}}1}}^{c_{t}^{k}}}$ is the relative pose of the target camera derived using the initialized continuous-time trajectory, and spatiotemporal parameters to be estimated. At this stage, all parameters within the estimator have been rigorously initialized.
III-D Continuous-Time Batch Optimization
Finally, a continuous-time-based bundle adjustment would be performed to refine all initialized parameters to the optimal states. The 3D-2D correspondences, organized from tracked grid patterns of the reference and target cameras, would be involved in constructing visual projection residuals for spatiotemporal optimization. The corresponding nonlinear least-squares problem can be expressed as follows:
$\small\hat{\mathcal{X}}\leftarrow\arg\min\sum_{k}^{\mathcal{P}_{\mathrm{ref}}}\sum_{j}^{\mathcal{G}^{k}_{\mathrm{ref}}}\rho\left(\left\|\boldsymbol{\mathrm{e}}_{\mathrm{ref}}^{k,j}\right\|^{2}\right)+\sum_{k}^{\mathcal{P}_{\mathrm{tar}}}\sum_{j}^{\mathcal{G}^{k}_{\mathrm{tar}}}\rho\left(\left\|\boldsymbol{\mathrm{e}}_{\mathrm{tar}}^{k,j}\right\|^{2}\right)$ (14)
with
$\boldsymbol{\mathrm{e}}_{\mathrm{ref}}^{k,j}$ $=\pi\left(\left({\hat{\boldsymbol{\mathrm{T}}}_{c_{r}}^{w}}(\tau^{k}_{j})\right)^{-1}\cdot\boldsymbol{\mathrm{p}}^{w}_{j},\mathcal{X}_{\mathrm{intr,ref}}\right)-\tilde{\boldsymbol{\mathrm{x}}}^{k}_{j}$ (15)
$\boldsymbol{\mathrm{e}}_{\mathrm{tar}}^{k,j}$ $=\pi\left(\left({\hat{\boldsymbol{\mathrm{T}}}_{c_{r}}^{w}}(\tau^{k}_{j}+{\hat{\tau}_{c_{t}}^{c_{r}}})\cdot{\hat{\boldsymbol{\mathrm{T}}}_{c_{t}}^{c_{r}}}\right)^{-1}\cdot\boldsymbol{\mathrm{p}}^{w}_{j},\mathcal{X}_{\mathrm{intr,tar}}\right)-\tilde{\boldsymbol{\mathrm{x}}}^{k}_{j}$
where $\boldsymbol{\mathrm{e}}_{\mathrm{ref}}^{k,j}$ and $\boldsymbol{\mathrm{e}}_{\mathrm{tar}}^{k,j}$ denote the projection residuals of 3D-2D pairs $\{\boldsymbol{\mathrm{p}}^{w}_{j},\tilde{\boldsymbol{\mathrm{x}}}^{k}_{j}\}$ of reference and target cameras, respectively; $\mathcal{X}_{\mathrm{intr,ref\mid tar}}$ are the intrinsics of two cameras; ${\hat{\boldsymbol{\mathrm{T}}}_{c_{r}}^{w}}(\tau)$ is the pose of $\underrightarrow{\mathcal{F}}_{c_{r}}$ in $\underrightarrow{\mathcal{F}}_{w}$ at time $\tau$, computed using the rotation and position B-splines; $\pi(\cdot)$ represents the visual projection function, which has been defined in (2); $\rho(\cdot)$ is the Huber loss function [40]. The nonlinear least-squares problems, i.e., (11), (12), and (14), would be solved using Ceres [41].
IV Real-World Experiment
To validate the feasibility of the proposed eKalibr-Stereo and evaluate its performance, comprehensive real-world experiments were conducted.
Figure 4: Stereo event camera rig (left subfigure) and three kinds of asymmetric circle grid patterns (right subfigures) utilized in real-world experiments.
IV-A Experiment Setup
Fig. 4 shows the self-assembled sensor rig for real-world experiments, consisting of two software-synchronized DAVIS346 event cameras (the resolution is 346$\times$260). We refer to the two event cameras as the left camera and the right camera for convenience in subsequent description and discussion. To ensure the comprehensiveness of the experiment, three different sizes of asymmetric circle grid patterns (3$\times$7, 4$\times$9, and 4$\times$11), as shown in Fig. 4, are used in real-world experiments. The radius rate and spacing for all grid boards are 2.5 and 50 mm, respectively. Altogether 5 sequences of 30-second data are collected for each grid board for Monte-Carlo evaluation.
To ensure the reliability of temporal calibration evaluation of eKalibr-Stereo in experiments, we manually shifted the timestamps of all events generated by the right camera by 50 ms after data acquisition, to simulate a stereo visual system with a time offset, i.e., we have $\tau_{\mathrm{right}}\leftarrow\tau_{\mathrm{right}}-\Delta\tau_{\mathrm{shift}}$ where $\Delta\tau_{\mathrm{shift}}=0.05$ sec. Therefore, the time offset of the right camera with respect to the left camera, i.e., ${\tau_{\mathrm{right}}^{\mathrm{left}}}$, is theoretically equal to 0.05 sec.
IV-B Evaluation and Comparison of Grid Pattern Tracking
Figure 5: Grid tracking performance of eKalibr-Stereo for three different sizes of ACircle grid boards. Based on the time distance threshold of neighbor patterns (set $\Delta\tau_{\mathrm{thd}}$ as 0.1 sec here) and the continuity threshold of grid tracking (set $\mathcal{N}_{\mathrm{thd}}$ as 50 here), tracked grids are segmented and corresponding piece-wise B-splines would be constructed. Details can be found in Section III-C1.
TABLE I: Evaluation and Comparison of Circle Gird Tracking
eKalibr-Stereo achieves the highest tracking success rate
Method Grid Type Cmp. Grid Incmp. Grid Total
eKalibr [21]    3×$\times 7$ 64.07 % 64.07 %
   4×$\times 9$ 57.43 % 57.43 %
   4×$\times 11$ 50.69 % 50.69 %
eKalibr-Stereo    3×$\times 7$ 64.07 % 27.63 % 91.69 %
   4×$\times 9$ 57.43 % 28.48 % 85.91 %
   4×$\times 11$ 50.69 % 39.17 % 89.86 %
    The tracking success rate is obtained by: the number of successful grid trackings divided by the total number of grids to track.
The performance of the grid tracking module described in Section III-B is first evaluated. Fig. 5 shows the tracking results of eKalibr-Stereo in three runs using different sizes of grid boards, where both complete and incomplete grid patterns were plotted. As can be seen, although eKalibr [21] is able to extract accurate complete grid patterns (those green ones), the corresponding continuity of tracking is poor. Building upon eKalibr, the proposed eKalibr-Stereo leverages the prior knowledge of motion continuity, to predict and track incomplete patterns using Lagrange polynomial, significantly improving the continuity of grid tracking. Based on these continuously tracked patterns, piece-wise B-splines can be effectively constructed (black lines), providing the necessary foundation for subsequent spatiotemporal optimization.
Table I further quantitatively summarizes the average rate of grid pattern tracking of eKalibr and eKalibr-Stereo in Monte-Carlo experiments. It can be seen that eKalibr achieves a relatively low pattern tracking rate, with an average of 57 %, which is primarily caused by high-dynamic motion required by spatiotemporal estimation (to ensure parameter observability). In contrast, eKalibr-Stereo can track incomplete grid patterns, significantly improving the tracking rate (with an average of 88 %, an increase of approximately 30 % compared with eKalibr) and corresponding tracking continuity.
IV-C Evaluation and Comparison of Calibration Performance
To comprehensively and quantitatively evaluate the spatiotemporal calibration performance of the proposed eKalibr-Stereo, we conducted real-world Monte-Carlo experiments. A total of three stereo visual calibration methods were incorporated in experiments for the evaluation and comparison of calibration results:
  1. 1.
    Frame-Based Stereo Calibration (DV Software [42]): The stereo frame-based (standard) visual extrinsic calibration toolkit provided by iniVation, i.e., the developer of DAVIS346 event camera employed in our real-world experiments. Since the DAVIS346 event camera supports standard frame output, stereo visual extrinsics can be accurately determined using the conventional frame-oriented calibration pipeline in DV Software. Therefore, the calibration results from DV Software can be treated as the reference. Note that DV Software only supports spatial (extrinsic) calibration, requiring the stereo event camera rig to be temporally synchronized.
  2. 2.
    Event-Based Image Reconstruction (E2VID [16]) & Kalibr [25]: To calibrate event cameras (such as DVXplorer [43]) that only support event output, a common approach is to ($i$) first use an event-based frame reconstruction algorithm to generate standard images, and then ($ii$) perform calibration using these reconstructed images in conventional calibrator. In our experiments, the event-based frame reconstruction method E2VID [16] and the well-known frame-based visual calibrator Kalibr [25] are utilized.
  3. 3.
    Event-Only eKalibr-Stereo: The proposed event-based stereo visual calibrator, supports both spatial (extrinsics) and temporal (time offset) determination.
TABLE II: Spatiotemporal Calibration Results in Monte-Carlo Experiments
eKalibr-Stereo achieves calibration accuracy and reliability comparable to conventional frame-based DV Software
Method Grid Type Extrinsic Temporal
Rotation (Euler angles, unit: degree) Translation (unit: cm) Time Offset (unit: ms)
Roll Pitch Yaw X Y Z Est. Ref.
DV Software [42]    3×$\times 7$0.47$\pm$0.01 0.70$\pm$0.02 0.38$\pm$0.02 12.07$\pm$0.01 -0.01$\pm$0.01 -0.38$\pm$0.04 50.00
   4×$\times 9$0.42$\pm$0.02 0.69$\pm$0.01 0.38$\pm$0.02 12.08$\pm$0.02 -0.01$\pm$0.01 -0.40$\pm$0.05 50.00
   4×$\times 11$0.45$\pm$0.01 0.68$\pm$0.02 0.39$\pm$0.01 12.07$\pm$0.02 -0.01$\pm$0.01 -0.39$\pm$0.05 50.00
E2VID [16] & Kalibr [25]    3×$\times 7$1.03$\pm$1.12 1.96$\pm$1.34 1.26$\pm$1.00 13.75$\pm$1.58 2.01$\pm$1.44 -2.23$\pm$1.39 57.37$\pm$4.48 50.00
   4×$\times 9$1.37$\pm$1.42 1.74$\pm$1.60 1.06$\pm$1.28 11.23$\pm$1.68 3.42$\pm$1.51 -1.72$\pm$1.25 55.92$\pm$4.03 50.00
   4×$\times 11$0.95$\pm$1.16 1.52$\pm$1.33 0.81$\pm$1.01 11.77$\pm$1.46 1.82$\pm$1.23 -1.55$\pm$1.19 54.57$\pm$3.99 50.00
eKalibr-Stereo    3×$\times 7$0.48$\pm$0.03 0.71$\pm$0.02 0.39$\pm$0.01 12.08$\pm$0.01 -0.01$\pm$0.01 -0.36$\pm$0.06 49.58$\pm$0.10 50.00
   4×$\times 9$0.36$\pm$0.01 0.67$\pm$0.02 0.38$\pm$0.01 12.07$\pm$0.01 -0.01$\pm$0.02 -0.42$\pm$0.06 49.98$\pm$0.06 50.00
   4×$\times 11$0.43$\pm$0.01 0.66$\pm$0.01 0.39$\pm$0.01 12.07$\pm$0.01 -0.01$\pm$0.01 -0.37$\pm$0.08 49.93$\pm$0.06 50.00
    All spatiotemporal parameters in this table, i.e., extrinsics and time offset, are those of the right camera with respect to the left camera.
    The value in each table cell is represented as (Estimate Mean) $\pm$ (STD).
Table II summarized final spatiotemporal calibration results of DV Software, E2VID & Kalibr, and eKalibr-Stereo in real-world Monte-Carlo experiments, showing the spatiotemporal estimates and corresponding standard deviations (STDs). As can be seen, the calibration method using E2VID and Kalibr achieved the poorest results with the largest STDs and biases compared with the frame-based DV Software. This is mainly due to the high noise in reconstructed image frames from E2VID (though reconstructed images are consistent on a macroscopic scale), which could reduce the extraction accuracy of the grid board in kalibr, further affecting the spatiotemporal determination. In comparison, eKalibr-Stereo is able to directly and accurately extract grid patterns from raw events, thereby achieving calibration results comparable to frame-based DV Software using continuous-time spatiotemporal optimization. Overall, eKalibr-Stereo can achieve an average STD of less than 0.05 degrees for extrinsic rotation determination, 0.05 cm for extrinsic translation, and 0.1 ms for time offset.
Figure 6: The distributions of projection errors after spatiotemporal optimization for two cameras in one solving. The subplots on the left correspond to the left camera, while the subplots on the right correspond to the right camera.
Fig. 6 plotted the distributions of projection errors of two cameras after spatiotemporal calibration using three different sizes of grid boards as visual targets in eKalibr-Stereo. It can be found that the projection errors follow a zero-mean normal distribution, indicating that the calibrated spatiotemporal parameters are well-estimated and unbiased. The average sigma of final projection errors is less than 0.1 pixels, indicating the high-accuracy spatiotemporal calibration eKalibr-Stereo capable of.
IV-D Evaluation of Computation Consumption
TABLE III: Computation Consumption in eKalibr-Stereo
Grid tracking consumed the majority of the processing time
Configuration OS Name Ubuntu 20.04.6 LTS 64-Bit
Processor 12th Gen Intel® Core™ i9
Graphics Mesa Intel® Graphics
Grid Type Computation Consumption (unit: minute)
Grid Tracking Optimization Total
   3×$\times 7$ 2.47 + 2.29 0.21 5.00
   4×$\times 9$ 2.65 + 2.64 0.25 5.54
   4×$\times 11$ 2.89 + 2.60 0.31 5.80
    The reported time represents the average time consumption across multiple (five) runs, each data sequence lasting 30 sec.
    The reported time in Grid Tracking means the average elapsed time for the left camera and right camera.
To evaluate the computation efficiency of eKalibr-Stereo, we recorded the runtime for each execution in the Monte-Carlo experiments and calculated the average time consumption. The corresponding results are summarized in Table III. It can be observed that eKalibr-Stereo takes approximately 5.5 minutes on average to calibrate a stereo event camera rig. In the calibration process, the majority of the time, approximately 90%, is spent on ACircle grid pattern extraction and tracking. The total time consumption increases with the size of the grid, which is reasonable.
V Conclusion
In this article, we present the proposed continuous-time-based spatiotemporal calibrator for event-based stereo visual systems, named eKalibe-Stereo, which is event-only and can accurately estimate both extrinsic and temporal parameters of the sensor suite. To improve the continuity of gird tracking, building upon eKalibr, an additional efficient procedure is designed in eKalibr-Stereo to track incomplete grid patterns. Based on tracked complete and incomplete grid patterns, a two-step initialization is first performed to recover the initial guesses of all parameters in the estimator, followed by a continuous-time batch optimization to refine all parameters to the optimal states. Extensive real-world experiments were conducted to evaluate the performance of the eKalibr-Stereo regarding grid tracking and spatiotemporal calibration. The results indicate that eKalibr-Stereo significantly improves the event-based grid tracking rate and could achieve spatiotemporal calibration accuracy comparable to frame-based stereo visual calibrators.
Acknowledgments
Shuolong Chen would like to express his gratitude to the School of Geodesy and Geomatics, Wuhan University, and the GREAT team for providing experimental equipment.
CRediT Authorship Contribution Statement
Shuolong Chen: Conceptualisation, Methodology, Software, Validation, Original Draft. Xingxing Li: Supervision, Funding Acquisition, Review and Editing. Liu Yuan: Data Curation, Review and Editing.